1.defenition 2.applications 3.uses 4.free lenth 5. Compressibility

1.defenition


2.applications


3.uses


4.free lenth


5.Compressibility

Grade:

1 Answers

Aravind Bommera
11 years ago

An Ultrasonic Interferometer is a simple and direct device to determine the ultrasonic velocity
in liquids with a high degree of accuracy.
The principle used in the measurement of velocity (V) is based on the accurate determination of
the wavelength (λ) in the medium. Ultrasonic waves of known frequency (f) are produced by a quartz
plate fixed at the bottom of the cell. These waves are reflected by a movable metallic plate kept
parallel to the quartz plate. If the separation between these two plates is exactly a whole multiple of
the sound wavelength, standing waves are formed in the medium. This acoustic resonance gives rise
to an electrical reaction on the generator driving the quartz plate and the anode current of the generator
becomes maximum.
If the distance is now increased or decreased and the variation is exactly one half wavelength
(λ/2) or multiple of it, anode current again becomes maximum. If the separation between quartz plate
and metallic plate is changed by d between two successive maximum anode current, then
d = λ/2...(1)
From the knowledge of wavelength (λ), the velocity (V) can be obtained by the relation :
V = λ × f =2df....(2)

 

Hit "Like" Please